# Perspective on Moreau–Yosida Regularization in Density-Functional Theory

### v1

@@ II.3 Moreau–Yosida regularization
How such a MY regularized functional looks like is shown with a one-dimensional example of a DFT functional in Figure 1 in Section III.1. This construction has a long list of favorable properties that we reproduces here, while the proofs can be found in standard textbooks [5; 7]. It is important to remember that we always demand at least Assumption 1, ensuring that $X$ and $X^{\ast}$ are both reflexive and strictly convex:
@@
    $f_{\varepsilon}$ is convex, finite, and Gâteaux differentiable. If in addition $X^{*}$ is uniformly convex (Assumption 2) then $f_{\varepsilon}$ is also Fréchet differentiable.
@@
    $\qquad\mathop{\mathrm{prox}}_{\varepsilon f}:X\to X,$ (16)
@@
    for each $u\in X^{*}$. By adding $\tfrac{\varepsilon}{2}\|u\|_{X^{*}}^{2}$ to $f^{*}(u)$ the $(f_{\varepsilon})^{*}$ automatically becomes a strictly convex function.
Note that $X$ strictly convex was already needed for the definition of the proximal map since it means that $x\mapsto\|x\|^{2}_{X}$ is strictly convex and thus the minimum in Eq. (16) is reached at a unique point. The reason that we can write $\mathop{\mathrm{prox}}_{\varepsilon f}$ instead of $\mathop{\mathrm{prox}}_{f}^{\varepsilon}$ (or similar) in order to include the regularization parameter $\varepsilon$ in the notation is that the definition of the proximal map could have also been written as
@@
In a Hilbert-space setting (Assumption 3), a few additional useful properties of MY regularization and the proximal map can be added [6]:
@@
    $f_{\varepsilon}$ is Fréchet differentiable and its functional derivative $\nabla f_{\varepsilon}$ is Lipschitz-continuous with constant $\varepsilon^{-1}$.
@@
    The proximal map is firmly nonexpansive, i.e., for all $x,y\in X$ it holds
@@
    $\qquad\geq\|\mathop{\mathrm{prox}}_{\varepsilon f}(x)-\mathop{\mathrm{prox}}_{\varepsilon f}(y)\|^{2}_{X}.$
@@
    $\partial(f_{\varepsilon})^{*}$ is strongly monotonous, i.e., for all $u,v\in X^{*}$ and $x\in\partial(f_{\varepsilon})^{*}(u)$, $y\in\partial(f_{\varepsilon})^{*}(v)$, it holds
@@
While the notion of firm nonexpansiveness itself can be generalized to Banach spaces [25], it does not necessarily hold for the proximal map anymore. Nevertheless, a result of uniform continuity can be achieved for the proximal map also in a certain Banach-space setting [26]. Combining v from the previous properties and the resolvent form for the proximal map yields
@@
which is called the Yosida approximation [6, Def. 23.1] of the multivalued mapping $\partial f$ in functional analysis. It was first published by Yosida around the same time as Moreau’s contributions to convex analysis and together they became the namesakes of the regularization method.
@@ II.4 Formal density-functional theory
The standard setting of DFT is the $N$-particle Hamiltonian with an external, scalar potential $v$ on $\mathbb{R}^{3}$, where we use atomic units,
@@
Here, we allow different choices of the interaction strength $\lambda$, which can range from the non-interacting case ($\lambda=0$) to the fully interacting case with a repulsive force ($\lambda=1$). For all choices of $\lambda\in\mathbb{R}$ and $v\in L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$, this Hamiltonian is self-adjoint and bounded below on its usual domain [27, Th. X.19]. It was already mentioned that the methods and results presented here readily apply to various settings outside of standard DFT, although subtleties around the involved spaces and definitions of the basic functionals have to be taken into account. A setting that also involves the vector potential is briefly discussed in Section VI.2.
@@
where the variation is either over all normalized wave functions or density matrices with finite kinetic energy and the required antisymmetry. The result, $E^{\lambda}(v)$, is then the ground-state energy of the system for a given external potential $v$ and chosen interaction strength. The principal idea of DFT is now to separate a density-dependent part inside this variation, which is universal in the sense that it does not depend on $v$ [28; 29; 30; 3]. An elegant way to do this is by Levy’s constrained search—that is, one varies over all wave functions or density matrices that yield a given one-particle density
@@
We write $\Psi\mapsto\rho$ for an $N$-particle wave function that has $\rho_{\Psi}=\rho$ and analogously for density matrices. We define the universal density functional by a constrained search over pure states and ensembles, respectively,
@@
where $\hat{H}^{\lambda}_{0}=\hat{T}+\lambda\hat{W}$. Since the constrained search over ensembles is over a larger set, we have $F_{\mathrm{DM}}^{\lambda}(\rho)\leq\tilde{F}^{\lambda}(\rho)$. The functional $F_{\mathrm{DM}}^{\lambda}(\rho)$ is convex since it is given as the infimum of a linear functional over a convex set, unlike $\tilde{F}^{\lambda}(\rho)$, which is therefore not necessarily convex whenever $\tilde{F}^{\lambda}(\rho)\neq F_{\mathrm{DM}}^{\lambda}(\rho)$ [31; 3], and explicit non-convex examples are known [32, Sec. VI.E].
The effective domain of a (convex) functional is the set of points where it is finite. For the constrained-search functionals $\tilde{F}^{\lambda}(\rho)$ and $F_{\mathrm{DM}}^{\lambda}(\rho)$, the effective domains agree and are also independent of $\lambda$ and we set
$D=\{x\in X\mid\tilde{F}^{\lambda}(x)<\infty\}=\{x\in X\mid F_{\mathrm{DM}}^{\lambda}(x)<\infty\},$ (29)
which is closed if the convex $F_{\mathrm{DM}}^{\lambda}(x)$ is lower semicontinuous. From the definition of the density $\rho_{\Psi}$ and the definition of the functionals it follows that any $\rho\in D$ must be non-negative and normalized to the number of particles. It is a fact that $D$ corresponds to the $N$-representable densities [3, Th. 1.2]. By definition we have $\tilde{F}^{\lambda}(x)=F_{\mathrm{DM}}^{\lambda}(x)=+\infty$ for $x\in X\setminus D$. For this reason, the usual convention will be to use the symbol $\rho$ for elements of $D$ that are ‘real’ densities, and use $x,y\in X$ otherwise.
The pure- and ensemble-state density functionals given in Eqs. (27) and (28), respectively, yield the same ground-state energy by the Hohenberg–Kohn variation principle,
@@
$F^{\lambda}(\rho)=(-E^{\lambda})^{*}(-\rho)\leq F_{\mathrm{DM}}^{\lambda}(\rho)$ (31)
as a new functional that is convex and lower semicontinuous by construction. We can then get $E^{\lambda}(v)$ from this functional as well,
@@
Now if the convex $F_{\mathrm{DM}}^{\lambda}(\rho)$ is also lower semicontinuous (Assumption 4 below), this implies
@@
Unlike convexity, the lower semicontinuity of $F_{\mathrm{DM}}^{\lambda}(\rho)$ does not follow automatically and needs to be checked for every choice of topology. In settings where $F_{\mathrm{DM}}^{\lambda}(\rho)$ is not lower-semicontinuous, the convex conjugate of $E^{\lambda}(v)$ does notably not retrieve the original constrained-search functional $F_{\mathrm{DM}}^{\lambda}(\rho)$ in Eq. (31). Thus it looses connection to the variational principle in Eq. (24) and consequently also to the equivalent ground-state problem involving the Schrödinger equation. This means that minimizers $\rho\in X$ in Eq. (30) can occur where the convex conjugate $F^{\lambda}(\rho)$ of $E^{\lambda}(v)$ is finite, yet the constrained-search functional $F_{\mathrm{DM}}^{\lambda}(\rho)=+\infty$. If, on the other hand, $v$ in $E^{\lambda}(v)$ supports a ground state $\Psi$, then $F_{\mathrm{DM}}^{\lambda}(\rho_{\Psi})$ is also finite. For this reason it needs to be stressed that albeit the requirement of lower semicontinuity for the ensemble constrained-search functional seems like a mathematical sophistry, it is a critical feature of the theory, and we will elaborate further on this on page 39. The convex $F^{\lambda}(\rho)=F_{\mathrm{DM}}^{\lambda}(\rho)$ on $X$ and the concave $E^{\lambda}(v)$ on $X^{*}$, as a conjugate pair, then serve as the backbone of a convex-analysis formulation of DFT.
@@
$X$ and $X^{*}$ must be such that the Hamiltonian $\hat{H}_{v}^{\lambda}$ is self-adjoint with a $v$-independent domain and bounded below, and the ensemble constrained-search functional $F_{\mathrm{DM}}^{\lambda}(\rho)$ is lower semicontinuous. This implies that the $N$-representability set $D$ is closed and $F_{\mathrm{DM}}^{\lambda}(\rho)=F^{\lambda}(\rho)$.
@@
In order to fulfill this assumption, we can think about Lieb’s choice $L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$ for the space of densities $X$, which indeed makes $F_{\mathrm{DM}}^{\lambda}(\rho)$ lower semicontinuous [3]. Yet, this space does not fulfill even Assumption 1, which is also not necessary for a consistent formulation of DFT. However, this assumption is needed for our later constructions involving MY regularization. Instead, we suggest to switch to a more general yet reflexive $X\supsetneq L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$. This, on the other hand, can break lower semicontinuity of the ensemble constrained-search functional if we switch to a coarser topology.
@@
The dual pairing between potentials and densities in $X$ suggests a matching space for potentials that consists of the topological dual $X^{*}$. For $L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$ the dual space is $L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$. If instead $X\supsetneq L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$, then $X^{*}\subsetneq L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$. We have to be careful that this smaller dual space still allows for a self-adjoint and bounded-below Hamiltonian on its usual domain, but indeed a result for such potentials is available [27, Th. X.19]. Interestingly, $L^{3/2}(\mathbb{R}^{3})$ is exactly the boundary case in terms of $L^{p}$ spaces and the Hamiltonian fails to be self-adjoint on its usual domain for $p<3/2$ when the potential can become too singular. This is the case for a potential $r^{-\alpha}$, $\alpha\geq 2$, where different self-adjoint realizations on the punctured domain $\mathbb{R}^{3}\setminus\{0\}$ become possible. To ensure a self-adjoint and bounded-below Hamiltonian on the usual domain, we thus need to limit the potential space to $X^{*}\subset L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$.
@@
At this point, another problem appears due to this limitation. In certain cases, this space will not accommodate a given physical setting—for example, it might exclude singular Coulomb potentials and thus molecular systems in a clamped nuclei or Born–Oppenheimer approximation. We can solve this problem by just including a fixed offset potential from a more general class,
$v_{\mathrm{o}}\in L^{2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})\subsetneq L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3}).$ (34)
@@
It can also be taken from any other potential class that still admits a self-adjoint and bounded-below Hamiltonian according to Assumption 4, including non-local potentials. In particular, $v_{\mathrm{o}}$ can fulfill the task of a binding potential for $N$ electrons. But one has to take careful note that such a potential can change the domain of the self-adjoint Hamiltonian if it is not from the class of Eq. (34) [27, Sec. X.2]. Instead of writing $E^{\lambda}(v_{\mathrm{o}}+v)$, we redefine the energy functional as
$E^{\lambda}(v)=\inf_{\Psi}\langle\Psi,\hat{H}^{\lambda}_{v_{\mathrm{o}}+v}\Psi\rangle,$ (35)
@@
including the fixed offset automatically. Note that this also affects the universal density functional as the corresponding Legendre–Fenchel transform that can be retrieved as
@@
While the inclusion of this offset potential does not change anything within the basic structure of DFT, which is the reason for not including it explicitly into our notation, it cannot be ignored when solving the underlying equations, and we will highlight its occurrence at several places in the following sections.
Whenever a minimizer can be found in the Hohenberg–Kohn variation principle of Eq. (30) with the functional $F_{\mathrm{DM}}^{\lambda}(\rho)$, we call this $\rho\in D$ a $v$-representable density. This means, in other words, that there is a $v\in X^{*}$ such that the (ensemble) ground-state solution $\hat{\Gamma}$ to the Schrödinger equation with Hamiltonian $\hat{H}^{\lambda}_{v_{\mathrm{o}}+v}$ yields $\hat{\Gamma}\mapsto\rho$ and ground-state energy $E^{\lambda}(v)=F_{\mathrm{DM}}^{\lambda}(\rho)+\langle v,\rho\rangle$. We next write out the relation in Eq. (31) as the Lieb variational principle,
@@
Since in general $F^{\lambda}(\rho)\leq F_{\mathrm{DM}}^{\lambda}(\rho)$ and since a $v$-representable $\rho$ is a minimizer in Eq. (30) and must fulfill $F_{\mathrm{DM}}^{\lambda}(\rho)=E^{\lambda}(v)-\langle v,\rho\rangle$, we see that $F^{\lambda}(\rho)=F_{\mathrm{DM}}^{\lambda}(\rho)$ always holds for $v$-representable densities, even when $F_{\mathrm{DM}}^{\lambda}(\rho)$ is not lower semicontinuous. In what follows we want to relate this property to the functional $F^{\lambda}(\rho)$ and its subdifferential.
The notions of sub- and superdifferentials serve an important purpose in the formulation of DFT. If there exists an optimizer $v\in X^{*}$ for a given $\rho\in X$ in Eq. (37), then the following are all equivalent for the lower-semicontinuous $F^{\lambda}(\rho)$,
@@
Yet, it is by no means secured that the set $\partial F^{\lambda}(\rho)$ is non-empty for all $\rho\in D$. But since $F^{\lambda}(\rho)$ is convex and lower semicontinuous by definition, it is known that its subdifferential is non-empty on a dense subset of the effective domain of the functional [5, Cor. 2.44]. But only if $F_{\mathrm{DM}}^{\lambda}(\rho)$ is lower semicontinuous too, we have $F_{\mathrm{DM}}^{\lambda}(\rho)=F^{\lambda}(\rho)$, and what was said above also applies to $F_{\mathrm{DM}}^{\lambda}(\rho)$. In addition, in this case, any $\rho\in\partial E^{\lambda}(v)$ is an optimizer in the second line of Eq. (30) and therefore also $\rho\in D$ . The convex set $\partial E^{\lambda}(v)\subseteq D$ then yields all possible ground-state densities for the potential $v\in X^{*}$. This $\partial E^{\lambda}(v)$ will be empty if $v$ does not support a ground state for interaction strength $\lambda$, includes a single element if $v$ supports a nondegenerate ground state, and contains multiple elements in case of degeneracy. We may now identify the points at which $F^{\lambda}(\rho)$ has a non-empty subdifferential with the ($v$-representable) ground-state densities of a Schrödinger equation. Hence, assuming a lower semi-continuous $F_{\mathrm{DM}}^{\lambda}(\rho)$, it also holds that the set of $v$-representable densities is dense in the set of $N$-representable densities. This way, it can be highlighted in a different way why the lower semicontinuity of the ensemble constrained-search functional is important. Else, with a $F_{\mathrm{DM}}^{\lambda}(\rho)$ that is not lower semicontinuous, it is possible to find $v\in-\partial F^{\lambda}(\rho)$ when at the same time $F_{\mathrm{DM}}^{\lambda}(\rho)=+\infty$, so this density has $\rho\in X\setminus D$ and would be clearly not $v$-representable. This is why Assumption 4 is important when dealing with the functionals $F_{\mathrm{DM}}^{\lambda}(\rho)$ $F^{\lambda}(\rho)$. See also the example given on page II.4.

### v2

@@ II.3 Moreau–Yosida regularization
[…]
Figure 1: Illustration of MY regularization of a convex $f(x)$ together with one of the parabolas that trace out $f_{\varepsilon}(x)$. The parabola touches $f_{\varepsilon}(x)$ at $x_{0}$ and has its vertex at the proximal point $\mathop{\mathrm{prox}}(x_{0})$. Further, three elements from the subdifferential at $x_{1}$ are depicted that are linear functions that fully lie below the convex functional.
MY regularization was originally introduced by Moreau [69] and can be linked to the previously discovered Yosida approximation of maximal monotone operators [110]. How such a MY regularized functional looks like is shown with one-dimensional examples in Figure 1 and Figure 2 in Section III.1. This construction has a long list of favorable properties that we reproduce here, while the proofs can be found in standard textbooks [8; 111]. It is important to remember that we always demand at least Assumption 1, ensuring that $X$ and $X^{\ast}$ are both reflexive and strictly convex:
@@
    $f_{\varepsilon}$ is convex, finite, and Gâteaux differentiable. If in addition $X^{*}$ is uniformly convex (Assumption 2), then $f_{\varepsilon}$ is also Fréchet differentiable.
@@
    $\qquad\mathop{\mathrm{prox}}_{\varepsilon f}\colon X\to X,$ (16)
@@
    for each $u\in X^{*}$. By adding $\tfrac{\varepsilon}{2}\|u\|_{X^{*}}^{2}$ to $f^{*}(u)$, the $(f_{\varepsilon})^{*}$ automatically becomes a strictly convex function.
Note that $X$ strictly convex was already needed for the definition of the proximal map since it means that $x\mapsto\|x\|^{2}_{X}$ is strictly convex and thus the minimum in Eq. (16) is reached at a unique point. The reason that we can write $\mathop{\mathrm{prox}}_{\varepsilon f}$ instead of $\mathop{\mathrm{prox}}_{f}^{\varepsilon}$ (or similar) in order to include the regularization parameter $\varepsilon$ in the notation, is that the definition of the proximal map could have also been written as
@@
In a Hilbert-space setting (Assumption 3), a few additional useful properties of the MY regularization and the proximal map can be added [9]:
@@
    $f_{\varepsilon}$ is Fréchet differentiable and its functional derivative $\nabla f_{\varepsilon}$ is Lipschitz continuous with constant $\varepsilon^{-1}$.
@@
    The proximal map is firmly nonexpansive, meaning that, for all $x,y\in X$, it holds
@@
    $\qquad\qquad\geq\|\mathop{\mathrm{prox}}_{\varepsilon f}(x)-\mathop{\mathrm{prox}}_{\varepsilon f}(y)\|^{2}_{X}.$
@@
    $\partial(f_{\varepsilon})^{*}$ is strongly monotonous, meaning that, for all $u,v\in X^{*}$ and $x\in\partial(f_{\varepsilon})^{*}(u)$, $y\in\partial(f_{\varepsilon})^{*}(v)$, it holds
@@
While the notion of firm nonexpansiveness itself can be generalized to Banach spaces [51], it does not necessarily hold for the proximal map anymore. Nevertheless, a result of uniform continuity can be achieved for the proximal map also in a certain Banach-space setting [3]. Combining v from the previous properties and the resolvent form for the proximal map, we obtain
@@
which exactly corresponds to the previously mentioned Yosida approximation [9, Def. 23.1] of the multivalued mapping $\partial f$. Together with Moreau’s contributions to convex analysis this justifies the naming of the regularization method.
@@ II.4 Formal density-functional theory
The standard setting of DFT is the $N$-particle Hamiltonian with an external, scalar potential $v$ on $\mathbb{R}^{3}$ (atomic units):
@@
Here, we allow different choices of the interaction strength $\lambda$, which can range from the non-interacting case ($\lambda=0$) to the fully interacting case with a repulsive force ($\lambda=1$). For all choices of $\lambda\in\mathbb{R}$ and $v\in L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$, this Hamiltonian is self-adjoint and bounded below on its usual form domain by the Kato–Lions–Lax-Milgram–Nelson (KLMN) theorem [83, Th. X.17 and Th. X.19]. It was already mentioned that the methods and results presented here readily apply to various settings outside of standard DFT, although subtleties around the involved spaces and definitions of the basic functionals have to be taken into account. A setting that also involves vector potentials is briefly discussed in Section VI.2.
@@
where the variation is over all normalized wave functions or density matrices with finite kinetic energy and the required antisymmetry. The result, $E^{\lambda}(v)$, is then the ground-state energy of the system for a given external potential $v$ and chosen interaction strength $\lambda$. The principal idea of DFT is now to separate off the only density-dependent part inside this variation, while the remaining part is universal in the sense that it does not depend on $v$ [47; 82; 61; 63]. An elegant way to achieve this is by Levy’s constrained search—that is, one varies over all wave functions $\Psi$ or density matrices that yield a given one-particle density that we define for wave functions and density matrices $\hat{\Gamma}=\sum_{j}p_{j}|\Psi_{j}\rangle\langle\Psi_{j}|$ as
@@
We define the universal density functional by a constrained search over pure states and ensembles, respectively,
@@
where $\hat{H}^{\lambda}_{0}=\hat{T}+\lambda\hat{W}$ and where we write $\Psi\mapsto\rho$ for an $N$-particle wave function with $\rho_{\Psi}=\rho$ and analogously for density matrices. Since the constrained search over ensembles includes the search over pure states, we have $F_{\mathrm{DM}}^{\lambda}(\rho)\leq\tilde{F}^{\lambda}(\rho)$. The functional $F_{\mathrm{DM}}^{\lambda}(\rho)$ is convex since it is given as the infimum of a linear functional over a convex set, unlike $\tilde{F}^{\lambda}(\rho)$, which is not necessarily convex. At non-pure-state $v$-representable $\rho$ it holds $\tilde{F}^{\lambda}(\rho)\neq F_{\mathrm{DM}}^{\lambda}(\rho)$ [62; 63], and explicit examples are known [79, Sec. VI.E].
The effective domain of a convex functional is the set of points where $f(x)<+\infty$. For the constrained-search functionals $\tilde{F}^{\lambda}(x)$ and $F_{\mathrm{DM}}^{\lambda}(x)$, the effective domains agree and are also independent of $\lambda$ and we set
$D$ $=\{x\in X\mid\tilde{F}^{\lambda}(x)<\infty\}$
$=\{x\in X\mid F_{\mathrm{DM}}^{\lambda}(x)<\infty\}.$ (29)
From the definition of the density $\rho_{\Psi}$ and the definition of the functionals, it follows that any $\rho\in D$ must be non-negative and normalized to the number of particles. It is a fact that $D$ corresponds to the $N$-representable densities [63, Th. 1.2]. Note that we generally use $\rho$ for elements in $D$ that are real densities, while we use $x$ for more general elements in $X$ that will later rise to prominence as ‘mixed densities’. By definition, we have $\tilde{F}^{\lambda}(x)=F_{\mathrm{DM}}^{\lambda}(x)=+\infty$ for $x\in X\setminus D$. For this reason, the usual convention will be to use the symbol $\rho$ for elements of $D$ that are ‘real’ densities, and to use $x,y\in X$ otherwise.
The pure- and ensemble-state density functionals given in Eqs. (27) and (28), respectively, yield the same ground-state energy in the Hohenberg–Kohn variational principle,
@@
$F^{\lambda}(\rho)=(-E^{\lambda})^{*}(-\rho)$ (31)
as a new functional that is convex and lower semicontinuous by construction. It needs to be stressed here that although $F^{\lambda}=(F_{\mathrm{DM}}^{\lambda})^{**}$ by the two previous equations, this does not necessarily mean that the two functionals agree, since $f^{**}=f$ only holds if the functional is convex and lower semicontinuous. It nevertheless generally holds that $F^{\lambda}(\rho)\leq F_{\mathrm{DM}}^{\lambda}(\rho)$. But in any case we can get $E^{\lambda}(v)$ back from $F^{\lambda}(\rho)$ by
@@
Now, if the convex $F_{\mathrm{DM}}^{\lambda}(\rho)$ is also lower semicontinuous (Assumption 4 below), this implies
@@
Unlike convexity, lower semicontinuity of $F_{\mathrm{DM}}^{\lambda}(\rho)$ must be checked for every choice of topology.
As said before, in settings where $F_{\mathrm{DM}}^{\lambda}(\rho)$ is not lower-semicontinuous, the convex conjugate of $E^{\lambda}(v)$ does notably not retrieve the original constrained-search functional $F_{\mathrm{DM}}^{\lambda}(\rho)$ in Eq. (31). Thus it looses connection to the variational principle in Eq. (24) and consequently also to the equivalent ground-state problem involving the Schrödinger equation. This means that minimizers $\rho\in X$ in Eq. (30) can occur where the convex conjugate $F^{\lambda}(\rho)$ of $E^{\lambda}(v)$ is finite, yet the constrained-search functional has $F_{\mathrm{DM}}^{\lambda}(\rho)=+\infty$. If, on the other hand, $v$ in $E^{\lambda}(v)$ supports a ground state $\Psi$, then $F_{\mathrm{DM}}^{\lambda}(\rho_{\Psi})$ is also finite. For this reason, it needs to be stressed that, albeit the requirement of lower semicontinuity for the ensemble constrained-search functional seems like a mathematical sophistry, it is a critical feature of the theory, and we will elaborate further on this on page II.4. The convex $F^{\lambda}(\rho)=F_{\mathrm{DM}}^{\lambda}(\rho)$ on $X$ and the concave $E^{\lambda}(v)$ on $X^{*}$, as a conjugate pair, then serve as the backbone of a convex-analysis formulation of DFT.
@@
$X$ and $X^{*}$ are such that the Hamiltonian $\hat{H}_{v}^{\lambda}$ is self-adjoint with a $v$-independent domain and bounded below, and the ensemble constrained-search functional $F_{\mathrm{DM}}^{\lambda}(\rho)$ is lower semicontinuous. This implies $F_{\mathrm{DM}}^{\lambda}(\rho)=F^{\lambda}(\rho)$.
@@
To fulfill this assumption, we can think about Lieb’s choice $L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$ for the space of densities $X$, which indeed makes $F_{\mathrm{DM}}^{\lambda}(\rho)$ lower semicontinuous [63]. Yet, this space does not fulfill Assumption 1, which is not necessary for a consistent formulation of DFT but needed for our later constructions involving MY regularization. Instead, we suggest to switch to a more general yet reflexive space $X\supsetneq L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$. This, on the other hand, can break lower semicontinuity of the ensemble constrained-search functional if we switch to a coarser topology.
@@
The dual pairing between potentials and densities in $X$ suggests a matching space for potentials that consists of the topological dual $X^{*}$. For $L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$, the dual space is $L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$. If instead $X\supsetneq L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$, then $X^{*}\subsetneq L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$. Interestingly, $L^{3/2}(\mathbb{R}^{3})$ is exactly the boundary case in terms of $L^{p}$ spaces and the Hamiltonian fails to be self-adjoint on its usual domain for $p<3/2$ when the potentials can become too singular. This is the case for a potential $r^{-\alpha}$, $\alpha\geq 2$, where different self-adjoint realizations on the punctured domain $\mathbb{R}^{3}\setminus\{0\}$ become possible. To ensure a self-adjoint and bounded-below Hamiltonian on the usual domain, we thus need to limit the potential space to $X^{*}\subseteq L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$.
@@
At this point, another problem appears due to this limitation. In certain cases, this space will not accommodate a given physical setting—for example, it might exclude singular Coulomb potentials and thus molecular systems in a clamped-nuclei or Born–Oppenheimer approximation. We can solve this problem by just including a fixed offset potential from our previous class
$v_{\mathrm{off}}\in L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3}).$ (34)
@@
It can also be taken from any other potential class that still admits a self-adjoint and bounded-below Hamiltonian according to Assumption 4, including non-local potentials. In particular, $v_{\mathrm{off}}$ can fulfill the task of a binding potential for $N$ electrons. But one has to take careful note that such a potential can change the domain of the self-adjoint Hamiltonian if it is not from the class of Eq. (34) [83, Sec. X.2]. Instead of writing $E^{\lambda}(v_{\mathrm{off}}+v)$, we redefine the energy functional as
$E^{\lambda}(v)=\inf_{\Psi}\langle\Psi,\hat{H}^{\lambda}_{v_{\mathrm{off}}+v}\Psi\rangle,$ (35)
@@
including the fixed offset automatically. Note that the offset potential then must also enter the definition of the universal density functional so that it remains compatible to $E^{\lambda}(v)$,
@@
While the inclusion of this offset potential does not change anything within the basic structure of DFT, which is the reason for not including it explicitly into the notation of the functionals, it cannot be ignored when solving the underlying equations, and we will highlight its occurrence at several places in the following sections.
Whenever a minimizer can be found in the Hohenberg–Kohn variation principle of Eq. (30) with the functional $F_{\mathrm{DM}}^{\lambda}(\rho)$, we call this $\rho\in D$ a $v$-representable density. This means, in other words, that there is a $v\in X^{*}$ such that the (ensemble) ground-state solution $\hat{\Gamma}$ to the Schrödinger equation with Hamiltonian $\hat{H}^{\lambda}_{v_{\mathrm{off}}+v}$ yields $\hat{\Gamma}\mapsto\rho$ and ground-state energy $E^{\lambda}(v)=F_{\mathrm{DM}}^{\lambda}(\rho)+\langle v,\rho\rangle$. We next write out the relation in Eq. (31) as the Lieb variational principle,
@@
Since in general $F^{\lambda}(\rho)\leq F_{\mathrm{DM}}^{\lambda}(\rho)$ and since a $v$-representable $\rho$ is a minimizer in Eq. (30) and must fulfill $F_{\mathrm{DM}}^{\lambda}(\rho)=E^{\lambda}(v)-\langle v,\rho\rangle$, we see that $F^{\lambda}(\rho)=F_{\mathrm{DM}}^{\lambda}(\rho)$ always holds for $v$-representable densities, even when $F_{\mathrm{DM}}^{\lambda}(\rho)$ is not lower semicontinuous. In what follows, we want to relate this property to the functional $F^{\lambda}(\rho)$ and its subdifferential.
The notions of sub- and superdifferentials serve an important purpose in the formulation of DFT. If there exists an optimizer $v\in X^{*}$ for a given $\rho\in X$ in Eq. (37), then the following statements are all equivalent for the lower-semicontinuous $F^{\lambda}(\rho)$,
@@
Yet, it is by no means guaranteed that the set $\partial F^{\lambda}(\rho)$ is non-empty for all $\rho\in D$. But since $F^{\lambda}(\rho)$ is convex and lower semicontinuous by definition, it is known that its subdifferential is non-empty on a dense subset of $D\subset X$, the effective domain of the functional [8, Cor. 2.44].
@@
If $F_{\mathrm{DM}}^{\lambda}(\rho)$ is lower semicontinuous, then $F_{\mathrm{DM}}^{\lambda}(\rho)=F^{\lambda}(\rho)$, and what was said above also applies to $F_{\mathrm{DM}}^{\lambda}(\rho)$. In addition, in this case, each $\rho\in\partial E^{\lambda}(v)$ is an optimizer in the second line of Eq. (30) and therefore $\rho\in D$ . The convex set $\partial E^{\lambda}(v)\subseteq D$ then yields all possible ground-state densities for the potential $v\in X^{*}$. This $\partial E^{\lambda}(v)$ will be empty if $v$ does not support a ground state for interaction strength $\lambda$, includes a single element if $v$ supports a nondegenerate ground state, and contains multiple elements in case of degeneracy. We may now identify the points at which $F^{\lambda}(\rho)$ has a non-empty subdifferential with the ($v$-representable) ground-state densities of a Schrödinger equation. Hence, assuming a lower semicontinuous $F_{\mathrm{DM}}^{\lambda}(\rho)$, it also holds that the set of $v$-representable densities is dense in the set of $N$-representable densities. In this way, it can be highlighted in a different manner why the lower semicontinuity of the ensemble constrained-search functional is important. Else, with a $F_{\mathrm{DM}}^{\lambda}(\rho)$ that is not lower semicontinuous, it is possible to find $v\in-\partial F^{\lambda}(\rho)$ when at the same time $F_{\mathrm{DM}}^{\lambda}(\rho)=+\infty$, so this density has $\rho\in X\setminus D$ and would clearly be not $v$-representable. This is why Assumption 4 is important when dealing with the functionals $F_{\mathrm{DM}}^{\lambda}(\rho)$ and $F^{\lambda}(\rho)$. See also the example given after the proposition below.